# Scalable Ensembling For Mitigating Reward Overoptimisation

### v1

@@ 1 Introduction
Modern language models have been ubiquitous in discussions of general purpose AI systems that can accomplish myriad tasks across many disciplines and with rapidly increasing capabilities OpenAI (2023); Manyika & Hsiao (2023); Touvron et al. (2023); Bommasani et al. (2022); Wei et al. (2022). However, alongside this increase in capabilities, there has been growing concern around the risks of such systems as they are not entirely interpretable and could be misused to cause substantial harm either maliciously or inadvertently (Hendrycks et al., 2023; Wang et al., 2023; Hendrycks et al., 2022).
@@
This reward over-optimization problem has been identified as a significant technical issue in scaling up learning from human feedback Ouyang et al. (2022); Casper et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Prior work has noted that because the reward models trained from user feedback are only a proxy for their underlying preferences, opting for a high reward can lead to performance degradation. Recent work has proposed mitigations either through ensembling or transforming the RL objective but these methods are either too computationally expensive or analytically intractable to be used with modern language models given limited compute (Coste et al., 2023; Eisenstein et al., 2023).
In our work, we propose a simple modification to a common strategy of ensembling: instead of maintaining multiple separate reward models, we opt for a shared backbone with different linear heads with the hypothesis that different initialization and training procedure will generate enough diversity. Our contributions show initial experiments validating the utility of such an approach in RLHF as it is just as performant as using a full ensemble for mitigating overoptimisation, while requiring less time for training during reward modeling, and less memory and time for PPO.
@@ 2 Related Work
Reinforcement Learning from Human Feedback (RLHF) RL is a powerful framework for learning diverse and performant policies that can tackle a wide array of tasks Sutton & Barto (2018); Abbeel & Ng (2004); Bellman (1957); Haarnoja et al. (2017); Drake (2005). Of particular recent interest are language models, which have shown impressive base capabilities in terms of instruction following. RLHF is a core component of the tuning processes for many of the currently highest-performing and most widely-used language models, improving their ability to follow instructions and align their responses with human preferences Stiennon et al. (2022); Christiano et al. (2023). However, tuning through RLHF can introduce a risk of overfitting to a proxy of true reward, since the reward model is learned. Prior work analyzes this phenomenon using a fixed “gold-standard” reward model in the place of humans, training proxy reward models from labels it provides (Gao et al., 2023).
Ensembles for Overoptimisation Recent work has shown multiple ways to mitigate this overoptimization issue, but the suggestions only apply during training or require expensive copies of multiple reward models for ensembling without deeper analysis as to why they mitigate this issue Coste et al. (2023). Related work further discovered that the ensembles are more effective when pretraining from scratch with separate seeds Eisenstein et al. (2023). Separately, other work has investigated uncertainty across ensemble members for RLHF showing that using the same backbone with different linear heads can lead to significant improvement in calibration, but that this weakly correlated with performance for summarization tasks Gleave & Irving (2022). More recent work investigates lightweight ensembles investigates a similar approach through uncertainty estimation. We draw on this literature to propose a multi-head reward modeling scheme with bayesian uncertainty estimation Zhang et al. (2024). We simply use a linear layer for each separate reward function in an ensemble and train the reward model by fine-tuning through the base encoder in the language model. Our approach allows utilizing the shared features from supervised fine-tuning but enabling the diversity of reward ensembles in a scalable manner with a simple min function.
@@ 4.1 Datasets and Motivation
We select the Alpaca Instructions dataset which emphasizes naturalistic, closed-form questions and answers, usually having a well-defined correct response, e.g., the number of planets in the solar system, but also included more open questions such as famous actors who started on broadway, why certain states are named, how to play kickball, etc. Taori et al. (2023). Given the most relevant prior work studying overoptimisation focused on this dataset, it is our primary dataset for training and evaluation (Coste et al., 2023).
@@ 4.2 Methodology and Tools
We follow the RLHF pipeline of supervised fine-tuning (SFT), reward learning, and proximal policy optimization (PPO). The experiments are implemented using the OPT model family Zhang et al. (2022). We opted for the Alpaca Farm codebase to serve as our framework (Dubois et al., 2023).
@@
We modify this pipeline by training the multi-head reward models with the base model initialized from SFT. When selecting the reward for a given sample, since we produce multiple predictions, we are presented with multiple choices for computing our final prediction Coste et al. (2023), but we simply take the minimum over the ensemble. A pessimistic estimator might reduce performance in principle, although for offline RL and RLHF preferences, it has been either conjectured or empirically validated that a minimum is optimal, and prior work demonstrated uncertainty penalties do not improve the efficiency of ensembles in preventing over-optimization over a simple min Zhu et al. (2023); Coste et al. (2023); Kumar et al. (2020); Chen et al. (2021). Our approach is shown in Figure 1, and all hyperparameters and further training details are in the Appendix.
@@ 4.3 Results
To test the efficacy of our approach, we run 1) Standard PPO 2) PPO with an ensemble of three reward models and 3) PPO with a multi-head reward model with three linear heads. We fix ensemble size at three as prior work found no gains when increasing to four or five members due to fixed compute, although we emphasize our method is amenable to a much larger number of ensemble members with minimal overhead Coste et al. (2023). We select the 1.3 B parameter OPT model as the proxy and the 6.7B variant as the gold model and run PPO for 15 epochs to extensively test overoptimisation. We also use the 1.3B model as the policy, after supervised fine-tuning.
We present our results in Figure 6. Given the difference in reward scales, and multiple works validating the benefit of using a full ensemble for overoptimisation we opt for two figures that demonstrate 1) the efficacy of the multi-head ensemble against standard PPO and 2) the efficacy of using a multi-head or full ensemble. Following prior work we also plot reward against KL divergence Gao et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Figure 6 clearly shows the replication of overoptimisation as we reproduce the concave-down nature of gold rewards under standard PPO and then summarily show how using a multi-head reward model bridges this gap. Figure 6 shows that the multi-head and full ensemble nature to similar gold performance, however we emphasize that our approach allows an ensemble with much larger reward models and following the prior work suggesting the full ensemble we train each member for three epochs to maximize performance whereas we find the multi-head approach needs only one epoch. We average over three seeds for each method and include further details and ablations in the appendix.
@@ 4.5 Calibration Analysis and Implications
We investigate the calibration of our multi-head reward models, as accurate representation of uncertainty is crucial in RLHF to prevent overoptimistic predictions Casper et al. (2023). Prior work found that ensembling improves calibration but is weakly related to model error when using a shared backbone with different linear heads, suggesting that separate reward models might be necessary for efficient ensembling Gleave & Irving (2022). However, we hypothesize that the difference in improvement is due to the diversity of tasks, as the AlpacaFarm dataset focuses on more open-ended questions compared to the summarization tasks in the prior work, which might increase model error due to epistemic uncertainty. We assess the calibration quality of the models using different ensemble objectives, such as mean and minimum reward estimation. The multi-head models, particularly with the minimum objective, demonstrate an enhanced ability to capture uncertainty in the reward signals without incurring significant computational overhead. While the minimum estimate is technically mis-calibrated, it leads to better than predicted performance at high levels of certainty. This suggests a regularization effect from the pessimistic estimator, as overoptimized models typically reach a local minima, explaining the ”U-shaped” curves that result as KL divergence from the base policy increases. Our findings align with previous work on pessimistic estimates in ensemble models, underscoring the benefits of a conservative approach in certain RL contexts. Furthermore, our experiments suggest that a smaller number of heads (3) might be optimal for mitigating overoptimisation, providing a valuable guideline for future research in ensemble-based models for RLHF. An increase in the number of heads beyond a certain threshold might introduce more noise than beneficial diversity, potentially leading to overfitting on dataset nuances, which is consistent with existing literature on the effectiveness of smaller ensembles in specific scenarios Galdran et al. (2023). See Figure 3 for details.
@@ 4.6 Conclusion
Our research contributes a novel method to improve the robustness of aligning language models by utilizing a robust pessimistic statistics while avoiding the cost of materializing a full ensemble. By addressing the challenges of overoptimization in a computationally efficient manner, we move a step closer to developing AI systems that can reliably align with human values and preferences. We plan to further investigate the effects of different datasets such as the more naturalistic Stanford Human Preferences Shah et al. (2019) and larger model sizes for future work.

### v2

@@ 1 Introduction
Modern language models have been ubiquitous in discussions of general-purpose AI systems that can accomplish myriad tasks across many disciplines and with rapidly increasing capabilities OpenAI (2023); Manyika & Hsiao (2023); Touvron et al. (2023); Bommasani et al. (2022); Wei et al. (2022). However, alongside this increase in capabilities, there has been growing concern around the risks of such systems as they are not entirely interpretable and could be misused to cause substantial harm either maliciously or inadvertently (Hendrycks et al., 2023; Wang et al., 2023; Hendrycks et al., 2022).
@@
This reward over-optimization problem has been identified as a significant technical issue in scaling up learning from human feedback Ouyang et al. (2022); Casper et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Prior work has noted that because the reward models trained from user feedback are only a proxy for their underlying preferences, opting for a high reward can lead to performance degradation. Recent work has proposed mitigations either through ensembling or transforming the RL objective. Still, these methods are either too computationally expensive or analytically intractable to be used with modern language models given limited compute (Coste et al., 2023; Eisenstein et al., 2023).
In our work, we propose a simple modification to a common strategy of ensembling: instead of maintaining multiple separate reward models, we opt for a shared backbone with different linear heads with the hypothesis that different initialization and training procedures will generate enough diversity. Our contributions show initial experiments validating the utility of such an approach in RLHF, as it is just as performant as using a full ensemble for mitigating overoptimization while requiring less time for training during reward modeling and less memory and time for PPO.
@@ 2 Related Work
Reinforcement Learning from Human Feedback (RLHF) RL is a powerful framework for learning diverse and performant policies that can tackle a wide array of tasks Sutton & Barto (2018); Abbeel & Ng (2004); Bellman (1957); Haarnoja et al. (2017); Drake (2005). Of particular recent interest are language models, which have shown impressive base capabilities in terms of instruction following. RLHF is a core component of the tuning processes for many of the currently highest-performing and most widely-used language models, improving their ability to follow instructions and align their responses with human preferences Stiennon et al. (2022); Christiano et al. (2023). However, tuning through RLHF can introduce a risk of overfitting to a proxy of true reward since the reward model is learned. Prior work analyzes this phenomenon using a fixed “gold-standard” reward model in place of humans, training proxy reward models from labels it provides (Gao et al., 2023).
Ensembles for Overoptimisation Recent work has shown multiple ways to mitigate this overoptimization issue, but the suggestions only apply during training or require expensive copies of multiple reward models for ensembling without deeper analysis as to why they mitigate this issue Coste et al. (2023). Related work further discovered that the ensembles are more effective when pretraining from scratch with separate seeds Eisenstein et al. (2023). Separately, other work has investigated uncertainty across ensemble members for RLHF, showing that using the same backbone with different linear heads can lead to significant improvement in calibration but that this weakly correlated with performance for summarization tasks Gleave & Irving (2022). Concurrent work investigates a similar approach through uncertainty estimation to propose a last layer reward modeling scheme with Bayesian uncertainty estimation Zhang et al. (2024b). Alternatively we use a linear layer for each separate reward function in an ensemble and train the reward model by fine-tuning through the base encoder in the language model. The recent work most similar to our approach investigates using either LoRA for fine-tuning a frozen transformer or similary fine-tuning linear heads with the same transformer backbone for a reward modeling ensemble Zhang et al. (2024a). Our approach differs in that we forgo using the mean or a least confidence bound estimate due to the observation that the min is better suited for minimizing epistemic uncertainty and is simpler to implement. This allows for utilizing the shared features from supervised fine-tuning while enabling the diversity of reward ensembles in a scalable manner with a simple minimum function.
@@ 4.1 Datasets and Motivation
We select the Alpaca Instructions dataset, which emphasizes naturalistic, closed-form questions and answers, usually having a well-defined correct response, e.g., the number of planets in the solar system, but also included more open questions such as famous actors who started on broadway, why certain states are named, how to play kickball, etc. Taori et al. (2023). Given the most relevant prior work studying overoptimisation focused on this dataset, it is our primary dataset for training and evaluation (Coste et al., 2023).
@@ 4.2 Methodology and Tools
We follow the RLHF pipeline of supervised fine-tuning (SFT), reward learning, and proximal policy optimization (PPO). The experiments use the OPT model family Zhang et al. (2022). We opted for the Alpaca Farm codebase as our framework (Dubois et al., 2023).
@@
We modify this pipeline by training the multi-head reward models with the base model initialized from SFT. When selecting the reward for a given sample, since we produce multiple predictions, we are presented with multiple choices for computing our final prediction Coste et al. (2023), but we simply take the minimum over the ensemble. A pessimistic estimator might reduce performance in principle. However, for offline RL and RLHF preferences, it has been either conjectured or empirically validated that a minimum is optimal, and prior work demonstrated uncertainty penalties do not improve the efficiency of ensembles in preventing over-optimization over a simple min Zhu et al. (2023); Coste et al. (2023); Kumar et al. (2020); Chen et al. (2021). Our approach is shown in Figure 1, and all hyperparameters and further training details are in the Appendix.
@@ 4.3 Results
To test the efficacy of our approach, we run 1) Standard PPO 2) PPO with an ensemble of three reward models and 3) PPO with a multi-head reward model with three linear heads. We fix ensemble size at three as prior work found no gains when increasing to four or five members due to fixed compute, although we emphasize our method is amenable to a much larger number of ensemble members with minimal overhead Coste et al. (2023). We select the 1.3 B parameter OPT model as the proxy and the 6.7B variant as the gold model and run PPO for 15 epochs to extensively test overoptimisation. We also use the 1.3B model as the policy after supervised fine-tuning.
We present our results in Figure 6. Given the difference in reward scales and multiple works validating the benefit of using a full ensemble for overoptimisation, we opt for two figures that demonstrate 1) the efficacy of the multi-head ensemble against standard PPO and 2) the efficacy of using a multi-head or full ensemble. Following prior work, we plot reward against KL divergence Gao et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Figure 6 clearly shows the replication of overoptimisation as we reproduce the concave-down nature of gold rewards under standard PPO and then summarily show how using a multi-head reward model bridges this gap. Figure 6 shows that the multi-head and full ensemble nature to similar gold performance however we emphasize that our approach allows an ensemble with much larger reward models and following the prior work suggesting the full ensemble we train each member for three epochs to maximize performance whereas we find the multi-head approach needs only one epoch. We average over three seeds for each method and include further details and ablations in the appendix.
@@ 4.5 Calibration Analysis and Implications
We investigate the calibration of our multi-head reward models, as an accurate representation of uncertainty is crucial in RLHF to prevent overoptimistic predictions Casper et al. (2023). Prior work found that ensembling improves calibration but is weakly related to model error when using a shared backbone with different linear heads, suggesting that separate reward models might be necessary for efficient ensembling Gleave & Irving (2022). However, we hypothesize that the difference in improvement is due to the diversity of tasks, as the AlpacaFarm dataset focuses on more open-ended questions compared to the summarization tasks in the prior work, which might increase model error due to epistemic uncertainty. We assess the calibration quality of the models using different ensemble objectives, such as mean and minimum reward estimation. The multi-head models, particularly with the minimum objective, demonstrate an enhanced ability to capture uncertainty in the reward signals without incurring significant computational overhead. While the minimum estimate is technically miscalibrated, it leads to better-than-predicted performance at high levels of certainty. This suggests a regularization effect from the pessimistic estimator, as overoptimized models typically reach a local minima, explaining the ”U-shaped” curves that result as KL divergence from the base policy increases. Our findings align with previous work on pessimistic estimates in ensemble models, underscoring the benefits of a conservative approach in certain RL contexts. Furthermore, our experiments suggest that a smaller number of heads (3) might be optimal for mitigating overoptimisation, providing a valuable guideline for future research in ensemble-based models for RLHF. An increase in the number of heads beyond a certain threshold might introduce more noise than beneficial diversity, potentially leading to overfitting on dataset nuances, which is consistent with existing literature on the effectiveness of smaller ensembles in specific scenarios Galdran et al. (2023). See Figure 3 for details.
@@ 4.6 Conclusion
Our research contributes a novel method to improve the robustness of aligning language models by utilizing a robust pessimistic statistics while avoiding the cost of materializing a full ensemble. By addressing the challenges of overoptimization computationally efficiently, we move a step closer to developing AI systems that can reliably align with human values and preferences. We plan to further investigate the effects of different datasets, such as the more naturalistic Stanford Human Preferences Shah et al. (2019) and larger model sizes for future work.
4.7 Acknowledgements
AMA acknowledges support from a Knight-Hennessy fellowship and a National Science Foundation fellowship. SK acknowledges support by NSF 2046795 and 2205329, NIFA award 2020-67021-32799, the Alfred P. Sloan Foundation, and Google Inc.